Joan can weed the garden in 40 min and her mother can weed it in 60 min. How long would it take them working together?
Answer by jorel1380(3719)

1/40+1/60=1/t
3/120+2/120=1/t
5/120=1/t
5t=120
t=120/5 minutes for Joan and her mother to weed the garden together..